# Video-as-Answer: Predict and Generate Next Video Event with Joint-GRPO

### v1

@@ 1 Introduction
While generative AI [48, 36, 21, 50, 38] has revolutionized text-based tasks in real-life domains like healthcare [13] and education [10], video generation models remain largely confined to entertainment [7, 52]. This is a missed opportunity, as video encapsulates rich, dynamic information about the physical world that text alone struggles to convey [28, 49].
Motivated by this observation, we pioneer Video-Next-Event Prediction (VNEP), a novel paradigm that uses video as the answer modality for Next-Event Prediction (NEP). While the established NEP task takes a video with a procedural or predictive question as input to predict the next event in text [16, 9, 31, 43], VNEP instead generates dynamic video responses. The shift from telling to showing enables VNEP to offer more intuitive and customized answers by leveraging video’s ability to convey spatial layout, motion, and temporal ordering, while adapting the demonstration to the user’s current state. As shown in Figure 1, the video answer guides the user through the remaining steps of the Windsor knot based on his current tie configuration (e.g., color, orientation and tightness), providing clarity and personalization that a generic text-only description cannot achieve.
However, VNEP introduces significant challenges that go beyond mere visual continuation. Unlike existing video extension task [55] which predicts frames based on spatiotemporal patterns (e.g., forecasting a ball’s trajectory), VNEP focus on event-conditioned reasoning. It requires a model to first comprehend the video and question, reason about the subsequent event from causal or procedural logic (e.g., inferring that adding soap is needed after observing dirty dishes being scrubbed with water), and then generate a video that is both visually coherent and semantically faithful to this inferred event. A straightforward solution is to employ a VLM for prediction followed by a VDM for generation. However, this cascaded pipeline suffers from a semantic-to-visual misalignment; the VLM’s textual output may be linguistically correct but visually unrealistic or unexecutable by the VDM, leading to semantically and visually divergent videos [44]. In contrast, unified models [23, 42, 34] attempt to align understanding and generation within a single model but face a capability trade-off, often excelling in one at the expense of the other and struggling to achieve optimal performance in both simultaneously [45]. Consequently, neither paradigm alone offers a satisfactory solution.
@@ 2 Related Work
Next-Event Prediction. The established NEP task requires predicting a future event given a video and a procedural or predictive question [9, 16, 39, 32]. Existing efforts predominantly address this as a textual NEP problem, focusing on generating descriptive answers. This line of work is supported by benchmarks like VLEP [17], MVP [33] and V1-33K [39], and leverages techniques ranging from event understanding [3] to multiscale temporal modeling and commonsense reasoning [4]. The recent rise of VLMs has further propelled this field, with methods that fine-tune on large-scale data [9] or utilize RL to elicit NEP capabilities from pre-trained models [39]. However, a fundamental limitation persists: the answer modality remains exclusively textual. Our work breaks from this paradigm by introducing VNEP, advancing next-event reasoning from textual description to dynamic video demonstration.
Group Relative Policy Optimization. Group Relative Policy Optimization (GRPO) was initially introduced by DeepSeek-Math [29] to enhance the reasoning capabilities of language models and align their outputs with human preferences. Its effectiveness has led to its adoption in video understanding and reasoning [18, 41, 11, 5, 12, 6], where it improves model performance on complex, open-ended queries. Beyond understanding, GRPO has also been applied to image and video generation [15, 44, 47, 25, 46]. In video generation, its primary role has been to enhance the alignment between the generated video and the text prompt [22, 46], as well as to improve consistency in reference-based generation tasks [30]. While these works apply GRPO to optimize a single model, our Joint-GRPO coordinates two models (a VLM for reasoning and a VDM for visualization) simultaneously, ensuring they are jointly aligned for the VNEP task.
@@ 3 VANS-Data-100K
Raw Data Collection. We collect data from two distinct sources to cover both procedural and predictive scenarios. For procedural data, we source high-resolution videos from COIN [51] and YouCook2 [54] to ensure clear visual demonstrations of step-by-step tasks. For predictive data, we gather videos from general-scene datasets [39, 2] and short-films [8], which are rich in narrative and causal dynamics.
@@
Clip Selection. We employ Gemini-2.5-Flash [35] as an automated quality filter to identify the optimal 3-5 second clip. For procedural data, it selects the clip that best aligns with the given caption. For predictive data, it first generates a detailed caption for each segment, ensuring the selected clip is both high-quality and semantically representative.
@@ 4 VANS
Figure 3 introduces the overall architecture of VANS. The input question is tokenized and fed into the VLM alongside high-level ViT visual features from the input video. We task the VLM with performing instruction-grounded reasoning to generate a textual caption describing the predicted next event, which serves as the semantic guide for the VDM. To ensure visual consistency, the VDM is conditioned on both the generated caption and low-level visual cues, which are extracted by tokenizing $n$ sampled input frames using a VAE [38]; these tokens are then concatenated into the VDM’s conditioning latent space. This enables fine-grained visual correspondence while generating novel scenes.
@@ 4.2 Joint-GRPO
  • $r_{t1}(s_{i},s_{\text{gt}})$ measures semantic similarity between generated and ground-truth captions using ROUGE-L [19].
  • $r_{v1}(v_{\text{out}}^{i},v_{\text{gt}})$ evaluates visual coherence of generated videos with ground-truth using CLIP Similarity [27].
@@
Through this two-stage optimization, the VLM and VDM co-evolve into a synergistic unit. The distinct, complementary roles of each reward component, along with the training reward curves, are presented in  B.
@@ 5.1 Settings
Benchmarks. We construct our evaluation benchmark by sampling 400 procedural and 400 predictive samples from our dataset with source videos from established benchmarks [54, 51, 2, 8] to ensure reliable ground-truth text & video answers. The evaluation set is strictly separated from the training data, with no video or question overlap.
@@
Baselines. As no existing methods are designed for VNEP, we adapt top-performing models from related fields to establish baselines. These include: (1) Video extension model Video-GPT [55]; (2) Cascaded pipelines formed by combining top-tire VLMs (Gemini-2.5-Flash [35], Qwen-2.5-VL-3B [1], and its NEP-finetuned version TEMPURA [9]) with VDMs (Wan-2.1-1.3B [38], FilmWeaver [24]); and (3) Unified model Omni-Video [34].
@@
Implementation Details. We initialize VANS with Qwen2.5-VL-3B as the VLM and Wan-2.1-1.3B as the VDM. For Video-GPT, we provide the input video and utilize its native capability for video continuation. For VANS and other baseline methods, we supply the input video and the corresponding question to perform NVEP. Implementation details of our VANS and baselines are presented in  C.
@@ 5.2 Main Results
Quantitative Comparisons. Table 1 shows that VANS performs favorably against all baselines. On procedural benchmarks, VANS (Joint-GRPO) achieves a ROUGE-L of 0.3631 and CLIP-V of 0.8021, outperforming the strongest cascaded baseline (Gemini-FilmWeaver at 0.2802 and 0.7102) and unified model (Omni-Video at 0.1075 and 0.6293). More importantly, Joint-GRPO brings a significant gain over the SFT version (e.g., ROUGE-L from 0.2812 to 0.3631 and CLIP-V from 0.7655 to 0.8021), demonstrating the effectiveness of our Joint-GRPO strategy. The Video Extension model Video-GPT yields the lowest CLIP-T (0.1997), as it generates frames without event reasoning. Please refer to  D for additional results.
@@ B.1 Reward Design
   YouCook2 [54] 9K
   COIN [51] 21K
@@
   Video-Holmes [8] 10K
   ActivityNet [2] 20K
   V1-33K [39] 10K
@@ C.1 Training of VANS
In the SFT stage, the VLM is trained for 10K steps using LoRA [14] (rank=8, alpha=32) with a learning rate of $5\times 10^{-5}$, while the VDM is fully fine-tuned for 20K steps across all DiT blocks with the same learning rate of $5\times 10^{-5}$.
For Joint-GRPO post-training, Stage 1 is optimized for 800 steps with a learning rate of $5\times 10^{-5}$. In Stage 2, to ensure the quality of anchor captions, we filter out those with ROUGE-L scores below 0.6 before proceeding with VDM adaptation. Stage 2 is then trained for 1K steps with the same learning rate of $5\times 10^{-5}$. We equip the VLM with LoRA (rank=8, alpha=32). For the VDM, we adopt the method from [22] to convert a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) to enable GRPO training. We set the KL coefficient $\beta=0.004$, clip range to $1\times 10^{-3}$, and sample group size to 8 per prompt.
@@ D.3 Generalization
Reasoning Image-to-Video Generation. VANS generalizes effectively to the reasoning image-to-video (I2V) task by treating a single image as a static video clip. This generalization capability is attributed to the model’s training on mixed datasets including Koala-36M [40] for I2V tasks. Figure 9 demonstrates examples from UI2V-Bench [53], when given an image of a banana and the instruction ”leave the banana for a week,” our model accurately predicts the temporal evolution, generating a video where the banana skin darkens. In contrast, other strong baselines struggle to capture this causal-physical transformation correctly. This demonstrates the robustness of our approach in understanding static visual contexts and reasoning about their potential dynamic futures.
@@ D.5 Video Results
All video results corresponding to the figures in this paper, along with additional examples, are provided in our project page: https://video-as-answer.github.io.

### v2

@@ 1 Introduction
While generative AI [yang2025qwen3, team2024gemini, liu2024deepseek, yang2024cogvideox, wan2025wan] has revolutionized text-based tasks in real-life domains like healthcare [goyal2024healai] and education [chu2025llm], video generation models remain largely confined to entertainment [cheng2025animegamer, yu2025gamefactory]. This is a missed opportunity, as video encapsulates rich, dynamic information about the physical world that text alone struggles to convey [searle1980minds, yang2024video].
Motivated by this observation, we pioneer Video-Next-Event Prediction (VNEP), a novel paradigm that uses video as the answer modality for Next-Event Prediction (NEP). While the established NEP task takes a video with a procedural or predictive question as input to predict the next event in text [lai2025cross, cheng2025TEMPURA, souvcek2025showhowto, wu2025stitch], VNEP instead generates dynamic video responses. The shift from telling to showing enables VNEP to offer more intuitive and customized answers by leveraging video’s ability to convey spatial layout, motion, and temporal ordering, while adapting the demonstration to the user’s current state. As shown in Figure 1, the video answer guides the user through the remaining steps of the Windsor knot based on his current tie configuration (e.g., color, orientation and tightness), providing clarity and personalization that a generic text-only description cannot achieve.
However, VNEP introduces significant challenges that go beyond mere visual continuation. Unlike existing video extension task [zhuang2025videogptclipdiffusion] which predicts frames based on spatiotemporal patterns (e.g., forecasting a ball’s trajectory), VNEP focus on event-conditioned reasoning. It requires a model to first comprehend the video and question, reason about the subsequent event from causal or procedural logic (e.g., inferring that adding soap is needed after observing dirty dishes being scrubbed with water), and then generate a video that is both visually coherent and semantically faithful to this inferred event. A straightforward solution is to employ a VLM for prediction followed by a VDM for generation. However, this cascaded pipeline suffers from a semantic-to-visual misalignment; the VLM’s textual output may be linguistically correct but visually unrealistic or unexecutable by the VDM, leading to semantically and visually divergent videos [xiao2025mindomni]. In contrast, unified models [luo2025univid, wei2025univideo, tan2025omni] attempt to align understanding and generation within a single model but face a capability trade-off, often excelling in one at the expense of the other and struggling to achieve optimal performance in both simultaneously [xiao2025haploomni]. Consequently, neither paradigm alone offers a satisfactory solution.
@@ 2 Related Work
Next-Event Prediction. The established NEP task requires predicting a future event given a video and a procedural or predictive question [cheng2025TEMPURA, lai2025cross, wang2025fostering, su2025eventformer]. Existing efforts predominantly address this as a textual NEP problem, focusing on generating descriptive answers. This line of work is supported by benchmarks like VLEP [lei2020more], MVP [tan2023multiscale] and V1-33K [wang2025fostering], and leverages techniques ranging from event understanding [chao2015hico] to multiscale temporal modeling and commonsense reasoning [chen2021joint]. The recent rise of VLMs has further propelled this field, with methods that fine-tune on large-scale data [cheng2025TEMPURA] or utilize RL to elicit NEP capabilities from pre-trained models [wang2025fostering]. However, a fundamental limitation persists: the answer modality remains exclusively textual. Our work breaks from this paradigm by introducing VNEP, advancing next-event reasoning from textual description to dynamic video demonstration.
Group Relative Policy Optimization. Group Relative Policy Optimization (GRPO) was initially introduced by DeepSeek-Math [shao2024deepseekmath] to enhance the reasoning capabilities of language models and align their outputs with human preferences. Its effectiveness has led to its adoption in video understanding and reasoning [li2025videochat, wang2025video, feng2025video, chen2025grpo, ge2025arc, chen2025exploring], where it improves model performance on complex, open-ended queries. Beyond understanding, GRPO has also been applied to image and video generation [jiang2025t2i, xiao2025mindomni, yan2025can, mi2025milr, xue2025dancegrpo]. In video generation, its primary role has been to enhance the alignment between the generated video and the text prompt [liu2025flow, xue2025dancegrpo], as well as to improve consistency in reference-based generation tasks [shen2025identity]. While these works apply GRPO to optimize a single model, our Joint-GRPO coordinates two models (a VLM for reasoning and a VDM for visualization) simultaneously, ensuring they are jointly aligned for the VNEP task.
@@ 3 VANS-Data-100K
Raw Data Collection. We collect data from two distinct sources to cover both procedural and predictive scenarios. For procedural data, we source high-resolution videos from COIN [coin] and YouCook2 [zhou2018towards] to ensure clear visual demonstrations of step-by-step tasks. For predictive data, we gather videos from general-scene datasets [wang2025fostering, caba2015activitynet] and short-films [cheng2025video], which are rich in narrative and causal dynamics.
@@
Clip Selection. We employ Gemini-2.5-Flash [team2023gemini] as an automated quality filter to identify the optimal 3-5 second clip. For procedural data, it selects the clip that best aligns with the given caption. For predictive data, it first generates a detailed caption for each segment, ensuring the selected clip is both high-quality and semantically representative.
@@ 4 VANS
Figure 3 introduces the overall architecture of VANS. The input question is tokenized and fed into the VLM alongside high-level ViT visual features from the input video. We task the VLM with performing instruction-grounded reasoning to generate a textual caption describing the predicted next event, which serves as the semantic guide for the VDM. To ensure visual consistency, the VDM is conditioned on both the generated caption and low-level visual cues, which are extracted by tokenizing $n$ sampled input frames using a VAE [wan2025wan]; these tokens are then concatenated into the VDM’s conditioning latent space. This enables fine-grained visual correspondence while generating novel scenes.
@@ 4.2 Joint-GRPO
  • $r_{t1}(s_{i},s_{\text{gt}})$ measures semantic similarity between generated and ground-truth captions using ROUGE-L [lin2004rouge].
  • $r_{v1}(v_{\text{out}}^{i},v_{\text{gt}})$ evaluates visual coherence of generated videos with ground-truth using CLIP Similarity [radford2021learning].
@@
Through this two-stage optimization, the VLM and VDM co-evolve into a synergistic unit. The distinct, complementary roles of each reward component, along with the training reward curves, are presented in Appendix B.
@@ 5.1 Settings
Benchmarks. We construct our evaluation benchmark by sampling 400 procedural and 400 predictive samples from our dataset with source videos from established benchmarks [zhou2018towards, coin, caba2015activitynet, cheng2025video] to ensure reliable ground-truth text & video answers. The evaluation set is strictly separated from the training data, with no video or question overlap.
@@
Baselines. As no existing methods are designed for VNEP, we adapt top-performing models from related fields to establish baselines. These include: (1) Video extension model Video-GPT [zhuang2025videogptclipdiffusion]; (2) Cascaded pipelines formed by combining top-tire VLMs (Gemini-2.5-Flash [team2023gemini], Qwen-2.5-VL-3B [bai2023qwen], and its NEP-finetuned version TEMPURA [cheng2025TEMPURA]) with VDMs (Wan-2.1-1.3B [wan2025wan], FilmWeaver [FilmWeaver]); and (3) Unified model Omni-Video [tan2025omni].
@@
Implementation Details. We initialize VANS with Qwen2.5-VL-3B as the VLM and Wan-2.1-1.3B as the VDM. For Video-GPT, we provide the input video and utilize its native capability for video continuation. For VANS and other baseline methods, we supply the input video and the corresponding question to perform NVEP. Implementation details of our VANS and baselines are presented in Appendix C.
@@ 5.2 Main Results
Quantitative Comparisons. Table 1 shows that VANS performs favorably against all baselines. On procedural benchmarks, VANS (Joint-GRPO) achieves a ROUGE-L of 0.3631 and CLIP-V of 0.8021, outperforming the strongest cascaded baseline (Gemini-FilmWeaver at 0.2802 and 0.7102) and unified model (Omni-Video at 0.1075 and 0.6293). More importantly, Joint-GRPO brings a significant gain over the SFT version (e.g., ROUGE-L from 0.2812 to 0.3631 and CLIP-V from 0.7655 to 0.8021), demonstrating the effectiveness of our Joint-GRPO strategy. The Video Extension model Video-GPT yields the lowest CLIP-T (0.1997), as it generates frames without event reasoning. Please refer to Appendix D for additional results.
@@ B.1 Reward Design
   YouCook2 [zhou2018towards] 9K
   COIN [coin] 21K
@@
   Video-Holmes [cheng2025video] 10K
   ActivityNet [caba2015activitynet] 20K
   V1-33K [wang2025fostering] 10K
@@ C.1 Training of VANS
In the SFT stage, the VLM is trained for 10K steps using LoRA [hu2022lora] (rank=8, alpha=32) with a learning rate of $5\times 10^{-5}$, while the VDM is fully fine-tuned for 20K steps across all DiT blocks with the same learning rate of $5\times 10^{-5}$.
For Joint-GRPO post-training, Stage 1 is optimized for 800 steps with a learning rate of $5\times 10^{-5}$. In Stage 2, to ensure the quality of anchor captions, we filter out those with ROUGE-L scores below 0.6 before proceeding with VDM adaptation. Stage 2 is then trained for 1K steps with the same learning rate of $5\times 10^{-5}$. We equip the VLM with LoRA (rank=8, alpha=32). For the VDM, we adopt the method from [liu2025flow] to convert a deterministic Ordinary Differential Equation (ODE) into an equivalent Stochastic Differential Equation (SDE) to enable GRPO training. We set the KL coefficient $\beta=0.004$, clip range to $1\times 10^{-3}$, and sample group size to 8 per prompt.
@@ D.3 Generalization
Reasoning Image-to-Video Generation. VANS generalizes effectively to the reasoning image-to-video (I2V) task by treating a single image as a static video clip. This generalization capability is attributed to the model’s training on mixed datasets including Koala-36M [wang2024koala36mlargescalevideodataset] for I2V tasks. Figure 9 demonstrates examples from UI2V-Bench [zhang2025ui2v], when given an image of a banana and the instruction ”leave the banana for a week,” our model accurately predicts the temporal evolution, generating a video where the banana skin darkens. In contrast, other strong baselines struggle to capture this causal-physical transformation correctly. This demonstrates the robustness of our approach in understanding static visual contexts and reasoning about their potential dynamic futures.
@@ D.5 Video Results
All video results corresponding to the figures in this paper, along with additional examples, are provided in supplementary.html.